# HySem: A context length optimized LLM pipeline for unstructured tabular extraction

### v1

@@ Abstract
Regulatory compliance reporting in the pharmaceutical industry relies on detailed tables, but these are often under-utilized beyond compliance due to their unstructured format and arbitrary content. Extracting and semantically representing tabular data is challenging due to diverse table presentations. Large Language Models (LLMs) demonstrate substantial potential for semantic representation, yet they encounter challenges related to accuracy and context size limitations, which are crucial considerations for the industry applications. We introduce HySem, a pipeline that employs a novel context length optimization technique to generate accurate semantic JSON representations from HTML tables. This approach utilizes a custom fine-tuned model specifically designed for cost- and privacy-sensitive small and medium pharmaceutical enterprises. Running on commodity hardware and leveraging open-source models, our auto-correcting agents rectify both syntax and semantic errors in LLM-generated content. HySem surpasses its peer open-source models in accuracy and provides competitive performance when benchmarked against OpenAI GPT-4o and effectively addresses context length limitations, which is a crucial factor for supporting larger tables.
HySem: A context length optimized LLM pipeline for unstructured tabular extraction
Narayanan PP1 Indian Institute of Science, Bangalore pnkrishnan27@gmail.com                        Anantharaman Palacode Narayana Iyer1 JNResearch Labs LLP, Bangalore ananth@jnresearch.com
@@ 1 Introduction
Tables are a fundamental method for presenting structured information across various industrial applications, including regulatory information in the pharmaceutical industry and financial reports for businesses. Some shared characteristics and inherent challenges in tables as mentioned by Fang et al. (2024) include Heterogeneity Borisov et al. (2023), Sparsity, Context-based interconnection Liu et al. (2023), Lack of prior knowledge Borisov et al. (2023), Borisov et al. (2024), etc. Recent advancements in Large Language Models (LLMs) have shown promise in addressing these challenges for various table related tasks. Some table related tasks and techniques include entity matching and data-imputation Li et al. (2023), Korini and Bizer (2023), tabular Q&A Chen et al. (2024), schema augmentation Narayan et al. (2022), Serialization Yin et al. (2020) Yan et al. (2024), Table Manipulation Zhang et al. (2024), Table Understanding Lobo et al. (2023), Prompt Engineering Zheng et al. (2023), Table RAG Sundar and Heck (2023). Role-play Zhao et al. (2023).
In typical companies, the vast knowledge contained within these tables often remains untapped for actionable insights due to the lack of effective tools and methodologies for extracting and interpreting the tabular data. For instance, documents related to regulatory compliance, such as Annual Product Quality Reviews (APQR), fulfill compliance requirements but are rarely utilized beyond their initial purpose. To leverage the abundant resources of knowledge present in tabular form and stored in various formats like PDF, Word or HTML it is essential to have an accurate extraction system that transforms the content in to a semantically meaningful format. Several works have investigated the LLM’s performance for different input tabular formats such as JSON formats Singha et al. (2023), HTML tables Sui et al. (2024), Markdown Narayan et al. (2022). Accurate semantic representations of tables in a structured form allows automated processing and that can significantly aid the downstream processes, particularly those that use Large Language Models (LLM) as a key component of their workflows. Extracting complex tabular data from semi-structured sources and providing a semantic representation is a hard problem due to the lack of any standards in tabular data representations. This results in a huge diversity and arbitrariness in the way real world tables are presented. The language understanding ability of LLMs and their proven results present an immense opportunity to develop semantics driven represen-
Figure 1: HySem Pipeline Architecture.
tations. However, LLM are prone to hallucinations Huang et al. (2023) Akhtar et al. (2023), the generated outputs may have accuracy challenges and the output formatting may even have syntax errors. In a number of cases, even state of the art LLMs are not very accurate in capturing the complete tabular information.
In this paper, we propose HySem, a context-optimized pipeline designed to generate accurate semantic JSON representations from HTML tables using a custom fine-tuned model. Our research goals were formulated in accordance with industry needs, particularly for small and medium pharmaceutical enterprises, where cost and privacy considerations are paramount.
1.1 Key Goals
  1. 1.
@@ 1.1 Key Goals
    On-Premise Model Deployment: Models should be run on-premise due to data security considerations.
  2. 2.
@@
    Commodity Hardware Compatibility: Models should operate on commodity hardware with widely available GPUs.
  3. 3.
    Model Size Limitation: The model size should be less than 10 billion parameters to ensure compatibility with GPUs supporting 16 GB RAM.
  4. 4.
@@
    Open-Source Software Stack: The software stack should leverage open-source tools and models.
  5. 5.
    Maximized Extraction Accuracy: Extraction accuracy should be maximized, as any inaccuracies would necessitate costly human correction efforts.
Guided by the above goals, our contributions include:
@@
  • Context Optimizer: A novel context optimizer that rewrites input text to significantly reduce the token count.
@@
  • Semantic Synthesizer: A custom fine-tuned model that generates accurate semantic representations of HTML tables.
  • Automated Syntax Correction: An automated correction system to address syntax errors.
@@
  • Evaluation Methodology: A novel evaluation methodology that measures the performance of our pipeline using both intrinsic and extrinsic metrics.
Input tokens Tokenizer Number of tokens
IRBESARTAN 75MG FILM COATED TABLETS ’IR’, ’B’, ’ES’, ’ART’, ’AN’, ’ ’, ’75’, ’MG’, ’ FIL’, ’M’, ’ CO’, ’ATED’, ’ TABLE’, ’TS’ 14
PANTOPRAZOLE 20 MG GASTRO - RESISTANT TABLETS ’P’, ’ANT’, ’OP’, ’RA’, ’Z’, ’OLE’, ’ ’, ’20’, ’ MG’, ’ G’, ’AST’, ’RO’, ’ -’, ’ RES’, ’IST’, ’ANT’, ’ TABLE’, ’TS’ 18
MPR:PNAHG:A:UK01/R6 ’MP’, ’R’, ’:P’, ’NA’, ’HG’, ’:A’, ’:’, ’UK’, ’01’, ’/R’, ’6’ 11
ML06:FPS:I:INAEG:AU01-002 ’ML’, ’06’, ’:’, ’FPS’, ’:I’, ’:’, ’INA’, ’EG’, ’:’, ’AU’, ’01’, ’-’, ’002’ 13
0619R11486 ’061’, ’9’, ’R’, ’114’, ’86’ 6
08/02/20 ’08’, ’/’, ’02’, ’/’, ’20’ 5
Table 1: Key terms used in a typical compliance document filed by the pharmaceutical industry and the corresponding tokens.
2 Method
@@
Our proposed LLM Pipeline, HySem, aims to address the challenge of converting unprocessed raw HTML tables into a semantic JSON data structure while addressing the limitations of LLMs concerning context length and inference time. By adopting a novel strategy for optimizing the number of tokens, we aim to meet industry standards for high accuracy and reduced inference time, particularly for regulatory compliance reporting. Our approach utilizes open-source LLMs with fewer than 10 billion parameters, suitable for running on commodity hardware.
@@
HySem achieves these requirements through a structured pipeline composed of three components: Context Optimizer Subsystem ($\mathcal{A}_{\text{CO}}$), Semantic Synthesizer ($\mathcal{A}_{\text{SS}}$), Syntax Validator ($\mathcal{A}_{\text{SV}}$). $\mathcal{A}_{\text{CO}}$ employs a novel pre-processing methodology for optimizing the context window utilized by HTML tables, enabling processing of large tables. $\mathcal{A}_{\text{SS}}$ utilises an LLM for transforming this optimized HTML table into semantic JSON. $\mathcal{A}_{\text{SV}}$ dissects the generated JSON for any syntax errors and outputs a syntactically-valid JSON.
Figure 2: Encoder subsystem of the Context Optimizer that rewrites the input tokens to an optimal set of tokens.
@@ 2.1 Context Optimizer Subsystem
The limited context size of large language models significantly constrains their ability to effectively process extensive tables. Although models with extended context lengths can alleviate this issue, they may sacrifice accuracy and increase processing times due to the quadratic complexity of self-attention mechanisms Tay et al. (2022) Vaswani et al. (2017). Recent advancements in Transformer architectures, such as Linformer Wang et al. (2020) and Flash Attention techniques Dao et al. (2022), have shown improvements in memory and time complexity. However, the necessity to reduce sequence length remains a fundamental challenge.
Figure 3: An example figure illustrating a table optimized by the Context Optimizer. (a) shows the HTML table in its original form without optimization and (b) shows the optimized HTML table. The cell content "MIC/MBC ($\mu$g/mL)" for example, represented in its original form in Figure 3(a), tokenizes into: [’MIC’, ’/M’, ’BC’, ’ (’, ’$\mu$’, ’ g’, ’/mL’, ’)’] containing 8 tokens. After optimization, the same cell content transforms into 2 tokens, represented in Figure 3(b) as: [’MIC’, ’/M’], resulting in a substantial reduction in token count.
Significant token inefficiency can occur when there is a misalignment between the tokenizer’s vocabulary and the domain-specific terminology in the input text. For example, the term "Amoxycillin," a common medication in pharmaceutical literature, is not part of the Llama 3 tokenizer’s vocabulary. As a result, it is represented by multiple tokens, thereby reducing token efficiency. This issue is further illustrated with the term "Potassium Clavulanate," which tokenizes as (’Pot’, ’assium’, ’ Cl’, ’v’, ’aul’, ’an’, ’ate’), resulting in seven tokens. Table 1. illustrates several key terms used in the pharmaceutical industry and their corresponding tokens obtained from Llama 3 tokenizer.
We introduce "Context Optimizer", a novel token alignment technique that performs dynamic token pruning to address these limitations and enhance context utilization. The Context Optimizer is implemented in 2 phases: Encoding phase and Decoding phase.
@@ 2.1.1 Encoding Phase
In the encoding phase, we begin with a preprocessing step where we remove tags and attributes from the HTML tables that do not contribute to semantic understanding, such as those meant for styling. This is followed by a minification step to further reduce the input size. By eliminating non-essential elements, we streamline the input data, making it more manageable for subsequent processing. A tokenizing step tokenizes the contents of each cell in the table. A cell may be empty in which case there are no tokens. A non-empty cell may transform to one or more tokens.
The goal of encoding is to arrive at representing the cell contents using a minimal set of tokens while ensuring a unique representation for each distinct cell content. The encoded inputs result in a valid HTML table with cell contents rewritten to require fewer tokens. A mapping table is maintained to track the original cell content and its optimized counterpart, facilitating later restoration. The encoded input is processed by the downstream modules in the HySem pipeline, including the Semantic Synthesizer and the Syntax Validator. Figure 3., illustrates a sample HTML table before and after optimization.
Collision Resolution: In the process of minimizing token sequences for efficient encoding of each cell content, it is possible for two distinct cell contents to map to some overlapping token sequences. To ensure each cell content maintains a unique representation, our method includes a collision resolution mechanism. The cell contents "4/8" and "4/4" in Figure 3(a) tokenize into ["4", "/", "8"] and ["4", "/", "4"], respectively, with the token sequence ["4", "/"] overlapping in both. After collision resolution, "4/8" is represented by the token sequence, ["4", "/"] while "4/4" transforms into ["4", "/", "4"], thus ensuring a unique representation as depicted in Figure 3(b).
@@ 2.1.2 Decoding Phase
In the decoding step, the output from the downstream processing is decoded to restore the original lexicon used in the table. This decoding step reverses the earlier encoding which is present in the model generated JSON, reconstructing the output to match the original input while retaining the benefits of the token optimization. The resulting JSON is both semantically accurate and more efficiently processed.
A key feature of our Context Optimizer is its dynamic nature, wherein the mapping of input words to an optimized token sequence and the corresponding decoding process are dependent solely on the current input, without relying on any static mapping.
@@ 2.2 Semantic Synthesizer
HTML tables are highly varied and arbitrary in nature, featuring elements such as row spans, column-spans, multi-level headers, nested tables and diverse data types. These intricacies make it challenging to accurately extract hierarchical and relational structures. Motivated by the LLM’s capability to comprehend deep semantic relationships, HySem adopts the open-sourced Meta-Llama-3-8B-Base model and fine-tunes it with a manually labeled dataset for transforming HTML tables into semantic JSON.
Concretely, the Semantic Synthesizer $\mathcal{A}_{\text{SS}}$ accepts a HTML table $\mathcal{H}_{\text{i}}$ pre-processed by the Context Optimizer, as input and produces JSON $\mathcal{J}_{\text{i}}$ as the output in the same encoded semantic space as the pre-processed input HTML table. Our initial experiments indicate that adopting Prompt Engineering to achieve this transformation results in the generation of less accurate JSON representations, as evidenced by our intrinsic and extrinsic evaluations. We found that the LLM is highly sensitive to prompts and struggles to effectively capture the wide variety present in these tables. We identified several common failure patterns present in the JSONs generated by the LLM, which are cataloged as Failure Modes in Table 2.
@@
Table 2: Common failure modes and their examples.
2.2.1 Dataset
PubTabNet: PubTabNet Smock et al. (2022) is a large-scale dataset for image-based table recognition, containing over 568,000 images of tables from scientific papers along with their corresponding HTML annotations.
FinTabNet: FinTabNet Zheng et al. (2021) is a dataset specifically designed for recognizing tables in financial documents, containing over 112,000 tables. Each table instance annotation includes fields such as the HTML structure and bounding box coordinates, similar to PubTabNet.
For our purposes, we extract the HTML from these sources and manually curate annotations for these tables. We filter tables that fit within the LLM’s context length (8k context window for Llama3). We randomly select and pre-process 1,364 HTML tables from both PubTabNet and FinTabNet using our Context Optimizer and manually curate corresponding JSON annotations. We split the data into 756 training samples and 608 testing samples. This dataset is used to train the Semantic Synthesizer. We chose these datasets as they are public and they are reasonable representations of the Pharmaceutical and Finance verticals respectively.
We also created hand labeled datasets to address the custom needs of the industry using the proprietary customer supplied data. These proprietary datasets are used to fine-tune models that use Semantic Synthesizer LLM as the base model.
2.2.2 Fine-tuning
We fine-tune the LLM in two stages. In the first stage, we utilize the open-sourced datasets discussed in the previous section, for fine-tuning our model. In the second stage, we manually curate JSON annotations for 130 randomly chosen tables from our proprietary APQR dataset and further fine-tune our model.
2.3 Syntax Validator

### v2

@@ Abstract
Regulatory compliance reporting in the pharmaceutical industry relies on detailed tables, but these are often under-utilized beyond compliance due to their unstructured format and arbitrary content. Extracting and semantically representing tabular data is challenging due to diverse table presentations. Large Language Models (LLMs) demonstrate substantial potential for semantic representation, yet they encounter challenges related to accuracy and context size limitations, which are crucial considerations for the industry applications. We introduce HySem, a pipeline that employs a novel context length optimization technique to generate accurate semantic JSON representations from HTML tables. This approach utilizes a custom fine-tuned model specifically designed for cost- and privacy-sensitive small and medium pharmaceutical enterprises. Running on commodity hardware and leveraging open-source models, HySem surpasses its peer open-source models in accuracy and provides competitive performance when benchmarked against OpenAI GPT-4o and effectively addresses context length limitations, which is a crucial factor for supporting larger tables.
@@ 1 Introduction
Data tables are essential for facilitating regulatory compliance and financial reporting across industries such as pharmaceuticals and finance. In the pharmaceutical sector, documents like Annual Product Quality Reviews (APQRs) often contain complex, ad-hoc structured tables that present significant challenges for integration into standard databases. While these tables hold valuable information, their unstructured format renders them non-queryable, hindering automated processing. Our objective is to convert these table-based assets, often stored in HTML format, into semantic JSON, which enables direct mapping of the generated JSON keys to the field names of the database schema, facilitating efficient integration and use in business analytics applications.
Processing real-world tables, particularly those prevalent in industries like pharmaceuticals, is a highly challenging task. Unlike structured databases with predictable schemas, pharmaceutical tables often have arbitrary placements of headers and data elements. Headers might appear mid-table, or there may be multiple, inconsistent levels of headers across different documents. Appendix A.2 Figure 3 illustrates a complex pharmaceutical data table.
The arbitrariness in table presentation renders rule-based approaches ineffective for transforming HTML tables into semantic JSON. Furthermore, developing and maintaining such algorithms is costly and does not scale efficiently to handle frequent changes in table formats. In contrast, large language models (LLMs) can be trained to recognize patterns and relationships within the data, offering a more robust, scalable and adaptable solution. However, LLMs often encounter performance challenges when handling complex tabular structures, particularly very long tables that contain multiple instances of pharma-specific terminology.
In this paper, we propose HySem, a pipeline designed to accurately transform HTML tables into semantic JSON representations (refer Appendix A.3 for detailed illustrations of semantic JSON). Central to our approach is a novel Context Optimiser, which employs a dynamic token pruning technique to rewrite the input HTML tables. This process reduces the token count while maintaining the semantic integrity of the data. As a result, HySem can efficiently manage large and complex tables, including those featuring specialized pharmaceutical terminology, without sacrificing performance.
Our research was motivated by the business needs of small and medium pharmaceutical enterprises, where cost and data privacy considerations are crucial. Specifically, our key goals are:
@@
  • On-Premise Model Deployment: Ensuring all models operate on-premise to maintain stringent data security standards.
@@
  • Compatibility with Commodity Hardware: Enabling models to run efficiently on commodity hardware with cost-effective, readily available GPUs.
  • Model Size Constraints: Limiting model size to under 10 billion parameters, making it suitable for GPUs with up to 16 GB of memory.
@@
  • Open-Source Software Stack: Leveraging open-source tools and models to minimize costs and foster accessibility.
Building on these foundational goals, we present the following contributions:
@@
  • Context Optimizer: A novel component that efficiently rewrites input data, significantly reducing token counts and enhancing processing speed.
@@
  • Semantic Synthesizer: A custom fine-tuned model designed to produce precise semantic representations from complex HTML tables.
  • Syntax Corrector: An agentic system that identifies and corrects syntax errors in the output JSON with minimal human oversight.
@@
  • Evaluation Methodology: A new framework that evaluates the performance of our approach using a combination of intrinsic and extrinsic metrics.
2 Methodology
@@
Our proposed LLM Pipeline, HySem, aims to address the challenge of converting unprocessed raw HTML tables into a semantic JSON data structure while addressing the limitations of LLMs concerning context length and inference time. By adopting a novel strategy for optimizing the number of tokens, we aim to meet the industry standards for high accuracy and reduced inference time, particularly for regulatory compliance reporting.
@@
HySem achieves these requirements through a structured pipeline composed of three components: Context Optimizer ($\mathcal{A}_{\text{CO}}$), Semantic Synthesizer ($\mathcal{A}_{\text{SS}}$), Syntax Corrector ($\mathcal{A}_{\text{SC}}$). $\mathcal{A}_{\text{CO}}$ employs a novel methodology for optimizing the context window utilized by HTML tables, enabling processing of large tables. $\mathcal{A}_{\text{SS}}$ transforms this optimized HTML table into semantic JSON. $\mathcal{A}_{\text{SC}}$ reviews the generated JSON for any syntax errors and outputs a syntactically-valid JSON. Figure 1 illustrates the Hysem architecture in detail.
Figure 1: Hysem Architecture diagram
@@ 2.1 Context Optimizer Subsystem
The limited context size of large language models significantly constrains their ability to effectively process large tables (Liu et al., 2024; Xu et al., 2024). Although models with extended context lengths can alleviate this issue, they may sacrifice accuracy and increase processing times due to the quadratic complexity of self-attention mechanisms (Sui et al., 2024; Tay et al., 2020; Vaswani, 2017). Recent advancements in Transformer architectures, such as Linformer (Wang et al., 2020) and Flash Attention techniques (Dao et al., 2022), have shown improvements in memory and time complexity. Nevertheless, the critical need to reduce token count persists, as it is essential for enabling faster inference times and optimizing the utilization of limited GPU memory, all without sacrificing model performance.
Our approach to optimizing context length utilization stems from the observation that significant token inefficiency arises when there is a mismatch between the tokenizer’s vocabulary and domain-specific terminology in the input text. For example, "Amoxycillin," a widely recognized pharmaceutical medication, is not present in the Llama 3 tokenizer’s vocabulary, leading to its representation by multiple tokens. In addition to domain-specific terminology, other data types commonly encountered in real-world tables, such as dates, string identifiers, and proper nouns, often fail to align with the tokenizer’s vocabulary, further increasing token count and exacerbating inefficiency.
We introduce Context Optimizer, a novel token alignment technique designed to minimize the impact of "token-vocabulary" misalignment. This approach optimizes token representation by aligning cell contents in a table with the tokenizer’s vocabulary, thereby reducing the number of tokens required to represent each cell.
The Context Optimizer operates in 2 phases: Encoding phase where we rewrite the cell contents into a more compact form and Decoding phase which restores the original contents.
@@ 2.1.1 Encoding Phase
In the encoding phase, our process is divided into two primary stages: standard pre-processing and a specialized token-based encoding method.
Pre-processing Steps
We begin with a set of conventional pre-processing steps. First, we remove tags and attributes from the HTML tables that do not contribute to semantic understanding, such as those meant for styling. After this, a minification step is applied, which strips away unnecessary white spaces, further optimizing the HTML table for encoding.
(a) HTML Table before optimization
(b) HTML Table after optimization
Figure 2: An illustration of a table optimized by the Context Optimizer is shown in 2(a), presenting the original HTML table without any optimization. 2(b) displays the same table after optimization. In these figures, tokens within each cell are highlighted with distinct colors to facilitate easy observation of the token count per cell. The tokenization is performed using the LLaMa 3 tokenizer.
Token-Based Encoding
Next, we apply our custom token-based encoding technique. The goal here is to represent the content of each cell with the minimum number of tokens while ensuring that each cell has a unique representation. The algorithm for this process is detailed in Algorithm 1.
Prior to encoding, we first sort the cells in ascending order based on the number of tokens they contain. This strategy allows us to resolve potential collisions more easily, as cells with fewer tokens are processed first. A collision occurs when two distinct cell contents map to overlapping token sequences.
Our encoding process incorporates several high-level heuristics to enhance efficiency and accuracy:
a. Single Token Preservation: If a cell’s content consists of a single token, it remains unchanged.
b. Multi-Token Optimization: For cells with multiple tokens, we aim to represent the content using only two tokens whenever possible. This approach addresses the initial attempt of encoding with a single token, which often compromised semantic richness and resulted in inaccurate JSON predictions.
c. Bracket Handling: We handle incomplete bracket sequences by checking if a token starts with an opening bracket (e.g., [, {) and lacks a matching closing bracket. In such cases, we concatenate subsequent tokens until the bracket is closed. This approach is crucial to prevent syntax errors in the generated JSON.
Figure 2 offers a detailed illustration of the HTML table before and after optimization. For example, in Figure 2(a) the cell content "Theme 1: Women’s knowledge and understanding of preeclampsia" is tokenized into 15 distinct tokens, represented by various colors. After optimization, the same cell content transforms into "Theme 1", consuming only 3 tokens as illustrated in Figure 2(b).
The overall objective is to use the fewest tokens possible while maintaining uniqueness across all cell contents. By treating each tokenized cell as a unit, we can reduce the total number of tokens significantly without losing semantic integrity. The encoded HTML table is processed by the downstream modules in the HySem pipeline, including the Semantic Synthesizer and the Syntax Corrector.
Algorithm 1 Token-Based Encoding Algorithm
Input: Table $T$ with $n$ cells $\{C_{1},C_{2},\dots,C_{n}\}$, each containing text.
Output: Encoded table $T^{\prime}=\{E_{1},E_{2},\dots,E_{n}\}$ with unique token representations.
Sort cells $\{C_{1},C_{2},\dots,C_{n}\}$ in ascending order based on the number of tokens in each cell.
for each cell $C_{i}\in T$ (after sorting) do
     Tokenize $C_{i}$: $T_{i}=\{t_{1},t_{2},\dots,t_{k}\}$ where $t_{j}$ are tokens of $C_{i}$.
     Apply heuristics to determine the initial encoded form $E_{i}$. $\triangleright$ Use predefined strategies
     if $\exists j<i$ such that $E_{i}=E_{j}$ then $\triangleright$ Check if a previous encoding $E_{j}$ is the same as $E_{i}$
         for $t_{k}\in\{t_{2},t_{3},\dots,t_{k}\}$ do
              $E_{i}=E_{i}\cup t_{k}$. $\triangleright$ Concatenate additional tokens
              if $E_{i}\neq E_{j}$ then
                  break. $\triangleright$ Stop once the encoding is unique
              end if
         end for
     end if
     Store the mapping $(C_{i},E_{i})$.
end for
return encoded table $T^{\prime}=\{E_{1},E_{2},\dots,E_{n}\}$.
@@ 2.1.2 Decoding Phase
In the decoding phase, the output generated by the LLM is decoded to restore the original lexicon used in the table. This decoding step reverses the earlier encoding which is present in the model generated JSON, reconstructing the content to accurately reflect the contents of the original HTML table while preserving the benefits of token optimization. The resulting JSON is both semantically accurate and more efficiently processed.
A key feature of our Context Optimizer is its dynamic nature, where the mapping of input words to optimized token sequences, as well as the corresponding decoding process, are entirely driven by the current input. This process operates without reliance on any pre-defined static mappings.
@@ 2.2 Semantic Synthesizer
Given the LLM’s capability to comprehend deep semantic relationships, HySem adopts the open-sourced Meta-Llama-3-8B-Base model and fine-tunes it with a manually labeled dataset for transforming HTML tables into semantic JSON.
Concretely, the Semantic Synthesizer $\mathcal{A}_{\text{SS}}$ accepts a HTML table $\mathcal{H}_{\text{i}}$ optimized by the Context Optimizer, as input and produces JSON $\mathcal{J}_{\text{i}}$ as the output in the same encoded semantic space as the optimized input HTML table. Our initial experiments indicate that adopting Prompt Engineering to achieve this transformation results in the generation of less accurate JSON representations, as evidenced by our Intrinsic and Extrinsic evaluations. We found that the LLM is highly sensitive to prompts and struggles to effectively capture the wide variety present in these tables. We identified several common failure patterns present in the JSONs generated by the LLM, which are cataloged as Semantic Failures in Table 1.
Dataset
PubTabNet: PubTabNet (Smock et al., 2022) is a large-scale dataset for image-based table recognition, containing over 568,000 images of tables from scientific papers along with their corresponding HTML annotations.
FinTabNet: FinTabNet (Zheng et al., 2021) is a dataset specifically designed for recognizing tables in financial documents, containing over 112,000 tables. Each table instance annotation includes fields such as the HTML structure and bounding box coordinates, similar to PubTabNet.
For our purposes, we extract the HTML from these sources and hand-label the corresponding JSON annotations for these tables. We filter tables that fit within the LLM’s context length (8k context window for Llama3) and select 1,364 HTML tables from both PubTabNet and FinTabNet. We split the data into 756 training samples and 608 testing samples. We chose these datasets as they are public and are reasonable representations of the Pharmaceutical and Finance verticals respectively.
We also created additional hand-labeled datasets to address the custom needs of the industry using the proprietary customer supplied data. These proprietary datasets are used to fine-tune models that use Semantic Synthesizer LLM as the base model.
@@
Table 1: Semantic Failures
Failure Mode Description
Missing List Enclosure The output is expected to be a list of dictionaries, but the LLM generates dictionaries without enclosing them in a list.
Unmatched Curly Braces The JSON output is missing one or more curly braces, resulting in an incomplete or unbalanced structure.
Missing Commas The JSON output is missing commas, leading to improperly formatted data.
Incorrect Placement of Quotes For long numeric strings, like 123,456,789, the output should enclose the entire string in quotes. Instead, the quotes are incorrectly inserted within the string, producing a format like 123,"456,789".
Table 2: Structural Failures
2.3 Syntax Corrector